OpenStudy (anonymous):

how do I graph each equation using its x and y intercept? 3x+2y=6

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

In order to find the x intersection, you have to set y=0 in the equation. In the same way to find the y-intercection, make x=0, that is: x-inter: 3x+2(0)=6 x=2 Answer 1 y-inter: 3(0)+2y=6 y=3 Answer 2

OpenStudy (anonymous):

Feed in values of x so that y = 0 then do the same for values of y so tht x = 0 You will have two coordinate points and you then join a line through them. Let x =0 then y = 3 Let y = 0 then x = 2 This gives the coordinates (0, 3)and (2, 0)

OpenStudy (anonymous):

Thank you so much I really appreciate it. Excellent job!!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

Do I do this problem the same way? y=x-2

OpenStudy (anonymous):

yeah exactly the same way. This method works for any linear equation.

OpenStudy (anonymous):

okay I think I got it enjoy your day

OpenStudy (anonymous):

np

OpenStudy (dboyette):

3x+2y=6 To find the x-intercept, substitute in 0 for y and solve for x. 3x+2(0)=6 Multiply 2 by each term inside the parentheses. 3x+0=6 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. 3x=6 Divide each term in the equation by 3. (3x)/(3)=(6)/(3) Simplify the left-hand side of the equation by canceling the common factors. x=(6)/(3) Simplify the right-hand side of the equation by simplifying each term. x=2 To find the y-intercept, substitute in 0 for x and solve for y. 3(0)+2y=6 Multiply 3 by each term inside the parentheses. 0+2y=6 Combine all similar expressions. 2y=6 Divide each term in the equation by 2. (2y)/(2)=(6)/(2) Simplify the left-hand side of the equation by canceling the common factors. y=(6)/(2) Solve the equation. y=3 These are the x and y intercepts of the equation 3x+2y=6. x=2, y=3